Cyber:

Type the correct answer in the box. Round your answer to the nearest whole number. Calculate the man’s mass. (Use PE = m × g × h, where g = 9.8 N/kg.) A man climbs a wall that has a height of 8.4 meters and gains a potential energy of 4,620 joules. His mass is about kilograms.

thematsercat20:

pot energy = mgh. 4620=m x10 x 8.4 (9.81->10) m = 4620/10x8.4 m=462/8.4 m=462/8 m=231/4 about 6kg, though a shade less ???

thematsercat20:

Can I get best response

GalaxyzStarz:

There is a formula to this. Plug the numbers into the equation

Cyber:

next question

Cyber:

Select the correct answer. An archer pulls back the string of a bow to release an arrow at a target. Which kind of potential energy is transformed to cause the motion of the arrow? A. chemical B. elastic C. gravitational D. magnetic

GalaxyzStarz:

A and D are defintely out

Cyber:

next Select all the correct answers. Which two objects have stored energy? a ball rolling on the ground a small rock sitting on top of a big rock a stretched rubber band a stone lying on the ground

Imagine:

Please place your question in another "Ask a Question" Box. It helps it get answered. Thanks!

Cyber:

were that at?